# Bergquist and Elmachtoub

### v4

@@ 3.1 Approximation of profit rate
Proof. Using Eq. (1), we first arrive at an upper bound for the revenue rate of the optimal policy
$\mathcal{R}(\pi^{*})=\sum_{i=0}^{\infty}\lambda_{i}^{*}p(\lambda_{i}^{*})\mathbb{P}_{i}(\pi^{*})\leq\tilde{\lambda}p(\tilde{\lambda})$ (6)
@@ 3.2 Approximation of revenue rate
We now state and prove Lemma 3.6 which ensures our constructed policy attains the advertised proportions of the optimal revenue.
@@ 3.3 Approximation of cost rate
which can be accomplished via the well-known Quartic formula due to Ferrari Cardano (1545).
For $\gamma\geq 3$, we proceed by showing that the objective is increasing in $\tilde{\lambda}$ over the interval $[0,1]$. Let $f_{\gamma}(\tilde{\lambda})$ denote the objective function for a fixed $\gamma$ and express it equivalently as
@@
We show the derivative $f^{\prime}_{\gamma}(\tilde{\lambda})$ is nonnegative for $0\leq\tilde{\lambda}\leq 1$ and $\gamma\geq 3$. We take the derivative and ignore the denominator by looking only at the signs
$\text{sgn}(f^{\prime}_{\gamma}(\tilde{\lambda}))$ $=\text{sgn}\left(\sum_{i=1}^{\gamma+2}\sum_{j=0}^{\gamma+1}j^{2}\tilde{\lambda}^{i+j-1}-\sum_{i=1}^{\gamma+2}\sum_{j=0}^{\gamma+1}ji\tilde{\lambda}^{i+j-1}\right)$
@@
We can use the above expression to write the coefficient of $\tilde{\lambda}^{k}$ for $12\leq k\leq\gamma+$ 1 as
@@
$\sum_{k=2}^{\gamma+1}\frac{(k-1)k(k+1)}{6}+\sum_{k=\gamma+2}^{2\gamma+2}\frac{(2\gamma-k+3)(4\gamma^{2}-4\gamma k+12\gamma+k^{2}-9k+8)}{6}=\frac{1}{12}(\gamma+1)(\gamma+2)^{2}(\gamma-3)$
@@ 3.4 A class of instances proving tightness
Consider an instance with linear demand (corresponding to a uniform valuation distribution which is regular), i.e. where $\lambda(p)=\max\{b-ap,0\}$ for some positive $a$ and $b$. Take $c=\mu=1$ and take $a<b<2a$, so that the expected value from accepting a customer when one is in service is always negative: the largest price a customer will pay is $b/a<2$, while the expected congestion penalty from accepting a customer while one is in service is $2$ (2 customers). Thus, the optimal policy is in fact a static policy with a cutoff point of 00. In this case, it is a straightforward single-variable optimization problem and we can arrive at the following expression for the optimal policy
@@
To now show the tightness of the objective value guarantee of Theorem 3.1, we write $b=\kappa a$ for $21<\kappa<$ 2 and take the limit as $a\rightarrow\infty$. In the limit, $\sqrt{\kappa a-a+1}-1\rightarrow\infty$, and thus $\tilde{\lambda}\rightarrow 1$. Then we have
@@ 4 Multi-server Static Pricing Guarantees
$g(\gamma,C)\mathcal{C}(\pi^{*})\geq\mathcal{C}(\pi^{\lambda,\gamma}).$
@@
We conjecture that the guarantees in Theorems 4.1 and 4.2 are tight, just as in the single‐server case. In fact, similar families of instances (e.g., those with linear demand, $\lambda=b-ap$, and appropriately chosen $a$ and $b$) appear to push each inequality in our proofs toward equality in multi‐server settings as well. However, identifying a closed‐form instance that rigorously demonstrates tightness is more challenging when $C>1$, due to the more complex state‐space and service parallelism. Our numerical experiments have not revealed any gap between these bounds and the observed performance, reinforcing our belief in tightness.
Figure 1: A graphical representation of the guarantees of Theorem 4.2
@@
$\tilde{\lambda}$ $=\sum_{i=0}^{\infty}\lambda_{i}\mathbb{P}_{i}(\pi)=\sum_{i=0}^{\infty}\lambda_{i}\frac{\prod_{k=0}^{i-1}\frac{\lambda_{k}}{\mu_{k+1}}}{1+\sum_{n=0}^{\infty}\prod_{k=0}^{n}\frac{\lambda_{k}}{\mu_{k+1}}}=\frac{\sum_{i=0}^{\infty}\frac{C\lambda_{i}}{C}\prod_{k=0}^{i-1}\frac{\lambda_{k}}{\mu_{k+1}}}{1+\sum_{n=0}^{\infty}\prod_{k=0}^{n}\frac{\lambda_{k}}{\mu_{k+1}}}\leq\frac{C\sum_{i=0}^{\infty}\prod_{k=0}^{i}\frac{\lambda_{k}}{\mu_{k+1}}}{1+\sum_{n=0}^{\infty}\prod_{k=0}^{n}\frac{\lambda_{k}}{\mu_{k+1}}}<C$
@@ 4.1 Approximation of profit rate
Proof. We compute the revenue rate of $\pi^{\tilde{\lambda},C-1}$ as
@@
$\Halmos\displaystyle\mathcal{Z}(\pi^{\tilde{\lambda},C-1})=\mathcal{R}(\pi^{\tilde{\lambda},C-1})-\mathcal{C}(\pi^{\tilde{\lambda},C-1})\geq\mathcal{Z}(\pi^{*})(1-\mathbb{P}_{C}(\pi^{\tilde{\lambda},C-1}))\geq\mathcal{Z}(\pi^{*})(1-\mathbb{P}_{C}(\pi^{C,C-1})).\Halmos$
4.2 Approximation of revenue rate
The following Lemma 4.6 proves the claimed revenue guarantees.
@@ 4.2 Approximation of revenue rate
4.3 Approximation of cost rate
The following Lemma 4.8 proves the claimed cost guarantees.
@@ 4.3 Approximation of cost rate
$g(\gamma,C)\mathcal{C}(\pi^{*})\geq\mathcal{C}(\pi^{\tilde{\lambda},C})$ (14)
@@
From Lemma 4.3, we know that that the maximum value of the coefficient of $\mathcal{C}(\pi^{*})$ is obtained for $\tilde{\lambda}\in[0,C]$. We can thus find the worst-case value and get a bound which does not depend upon $\tilde{\lambda}$. Indeed, letting $f(\gamma,C,\tilde{\lambda})=\frac{\sum_{i=1}^{\gamma+1}i\mathbb{P}_{i}(\pi^{\tilde{\lambda},\gamma})}{\tilde{\lambda}}$, we can write $g(\gamma,C)=\max_{\tilde{\lambda}\in[0,C]}f(\gamma,C,\tilde{\lambda})$ so that we have
@@ 5 Sojourn Time Penalty
where the second equality uses Little’s Law. Let $\pi^{*}_{s}$ denote the optimal policy for $\mathcal{Z}_{s}(\cdot)$ which has arrival rates $\lambda^{*}_{s,i}$. We use the notation $\tilde{\lambda}_{s}$ to denote the expected arrival rate under $\pi^{*}_{s}$, i.e, $\tilde{\lambda}_{s}=\sum_{i=0}^{\infty}\lambda^{*}_{s,i}\mathbb{P}_{i}(\pi^{*}_{s})$. In this case, objective value guarantees on $\mathcal{Z}_{s}(\cdot)$ for the class of static policies which pick the arrival rate $\tilde{\lambda}_{s}$ are no longer possible. This fact is verified with an instance from the class described in Section 3.4. If we take an instance with linear demand ($\lambda=b-ap$) with $b=6000$ and $a=5000$, we can observe that $\mathcal{Z}_{s}(\pi^{*}_{s})\approx 0.17>0$, whereas $\mathcal{Z}_{s}(\pi^{\tilde{\lambda}_{s},0})\approx-0.4<0$. For these instances, taking a cutoff higher than 00 degrades performance further. Thus no result of the flavor of Theorem 4.1 is possible, and hence we must resort to providing bi-criteria bounds on revenue and congestion in the spirit of Theorem 4.2.
[…]
@@
$E[W(\pi^{\tilde{\lambda}_{s},\gamma})]\leq E[W(\pi^{*}_{s})]E[W(\pi^{\tilde{\lambda},\gamma})]\leq E[W(\pi^{*}_{s})]E[W(\pi^{C,\gamma})]=E[W(\pi^{*}_{s})]\frac{\sum_{i=1}^{\gamma+1}i\mathbb{P}_{i}(\pi^{C,\gamma})}{C(1-\mathbb{P}_{\gamma+1}(\pi^{C,\gamma}))}$
@@ 6 Numerical Experiments
In this section, we present the results of numerical experiments on our static policies. Specifically, we consider linear ($\lambda=b-ap$), exponential ($\lambda=be^{-ap}$), and logistic ($\lambda=\frac{b\,(1+e^{-ap_{0}})}{1+e^{a(p-p_{0})}}$) demand functions. We perform experiments with $1,3,5,$ and $10$‐server systems. Though these systems may be relatively small in the context of applications like food delivery and cloud computing, we will see that already with $C=10$ the performance is near-optimal, so further experiments would not be illustrative. For each number of servers $C\in\{1,3,5,10\}$, we sample $a$ uniformly from $[0.1,5]$, $b$ uniformly from [.$[.5,10]$, and $p_{0}$ uniformly from $[0,20]$. Under linear demand, when $a>b$, the optimal policy does not sell and thus achieves an objective value of 00. Though this is also attainable by a static policy, for simplicity we reject these instances and resample.
@@
    Truncation: For an $M$‐state queue, we choose $M$ large enough so that the stationary probability of being at state $M$ is below a small threshold (e.g., 1%). This ensures we capture essentially all of the probability mass.
@@ 7 Conclusion
There are many possible avenues to extend our results. One promising direction of interest would be to relax the assumption that the service times are exponential. With general i.i.d. service times, the optimal pricing policy now depends not only on the number in system but also on the remaining service times of the customers in service. There are many subtleties that must be accounted for to prove similar results, but at the heart of our argument are two very robust tools: Little’s Law and Jensen’s inequality. Another interesting direction is to consider static pricing in queueing networks that represent vehicle sharing systems, where fair policies are of general concern (Elmachtoub and Kim 2024).

### v5

@@ 3.1 Approximation of profit rate
Proof of Theorem 3.1. Using Eq. (1), we first arrive at an upper bound for the revenue rate of the optimal policy
[…]
@@ 3.2 Approximation of revenue rate
We now state and prove Lemma 3.6 which ensures our constructed policy attains the revenue guarantee in Theorem 3.2.
@@ 3.3 Approximation of cost rate
which can be accomplished via the well-known Quartic formula due to Ferrari ArsMagna.
[…]
@@
[…]
$)\displaystyle\text{sgn}(f^{\prime}_{\gamma}(\tilde{\lambda}$)) $=\text{sgn}\left(\sum_{i=1}^{\gamma+2}\sum_{j=0}^{\gamma+1}j^{2}\tilde{\lambda}^{i+j-1}-\sum_{i=1}^{\gamma+2}\sum_{j=0}^{\gamma+1}ji\tilde{\lambda}^{i+j-1}\right)$
@@
We can use the above expression to write the coefficient of $\tilde{\lambda}^{k}$ for $2\leq k\leq\gamma+1$ as
@@
$\sum_{k=2}^{\gamma+1}\frac{(k-1)k(k+1)}{6}$ $+\sum_{k=\gamma+2}^{2\gamma+2}\frac{(2\gamma-k+3)(4\gamma^{2}-4\gamma k+12\gamma+k^{2}-9k+8)}{6}$
$=\frac{1}{12}(\gamma+1)(\gamma+2)^{2}(\gamma-3).$
@@ 3.4 A class of instances proving tightness
Consider an instance with linear demand (corresponding to a uniform valuation distribution which is regular), i.e. where $\lambda(p)=\max\{b-ap,0\}$ for some positive $a$ and $b$. Take $c=\mu=1$ and take $a<b<2a$, so that the expected value from accepting a customer when one is in service is always negative: the largest price a customer will pay is $b/a<2$, while the expected congestion penalty from accepting a customer while one is in service is $2$ (2 customers). Thus, the optimal policy is in fact a static policy with a cutoff point of 0. In this case, it is a straightforward single-variable optimization problem and we can arrive at the following expression for the optimal policy
@@
To now show the tightness of the objective value guarantee of Theorem 3.1, we write $b=\kappa a$ for $1<\kappa<2$ and take the limit as $a\rightarrow\infty$. In the limit, $\sqrt{\kappa a-a+1}-1\rightarrow\infty$, and thus $\tilde{\lambda}\rightarrow 1$. Then we have
@@ 4 Multi-server Static Pricing Guarantees
$g(\gamma,C)\mathcal{C}(\pi^{*})\geq\mathcal{C}(\pi^{\tilde{\lambda},\gamma}).$
@@
We conjecture that the guarantees in Theorems 4.1 and 4.2 are tight, just as in the single‐server case. In fact, similar families of instances (e.g., those with linear demand, $\lambda=b-ap$, and appropriately chosen $a$ and $b$) appear to push each inequality in our proofs toward equality in multi‐server settings as well. However, identifying a closed‐form instance that rigorously demonstrates tightness is more challenging when $C>1$, due to the more complex state‐space and service parallelism.
4.1 Nearly tight examples
We now provide some empirical support for our conjectured tightness of the profit, revenue, and cost guarantees. First, consider the profit guarantee of Theorem 4.1. We take the case $C=5$ with linear demand $\lambda=b-ap$, where a=10,000,000a=10{,}000{,}000 and $b=\left(1+\tfrac{0.999}{C}\right)a$. In this instance, the optimal profit is $0.996264$, while the profit under the static cutoff policy arrival rate $\tilde{\lambda}$ and $\gamma=4$ is $0.7136287$, yielding a ratio of $0.7163$. This is quite close to the theoretical guarantee of $0.715$ furnished by Theorem 4.1, lending credence to our belief that the bound is sharp.
Next, we examine the revenue guarantee of Theorem 4.2 in the same instance. The optimal revenue is $5.9835687$, while the static policy with $\gamma=6$ achieves $4.9073757$, for a ratio of $0.8196$. This value is very close to the theoretical guarantee of $0.8185$. As $a\to\infty$, we expect this ratio to converge even more tightly to the bound. This experiment demonstrates that the simple linear-demand construction mirrors the single-server worst case and confirms the essential sharpness of our revenue guarantee.
Finally, we illustrate the tightness of the cost guarantee of Theorem 4.2. Here we consider $C=5$ with $a=66.7$, $b=73.4$, and cutoff $\gamma=6$. In this case, the optimal cost is $2.59288$, while the static policy yields $2.6188$. This gives a ratio of 1.0100285516, nearly the same as the theoretical guarantee $g(6,5)=1.0100285581$. Together, these experiments lend credence to our conjecture profit, revenue, and cost guarantees are tight in the multi-server case as well.
4.2 Proofs of Theorems 4.1 and 4.2
@@ 4.1 Nearly tight examples
$\tilde{\lambda}$ $=\sum_{i=0}^{\infty}\lambda_{i}\mathbb{P}_{i}(\pi)$
$=\sum_{i=0}^{\infty}\lambda_{i}\frac{\prod_{k=0}^{i-1}\tfrac{\lambda_{k}}{\mu_{k+1}}}{1+\sum_{n=0}^{\infty}\prod_{k=0}^{n}\tfrac{\lambda_{k}}{\mu_{k+1}}}$
$=\frac{\sum_{i=0}^{\infty}\tfrac{C\lambda_{i}}{C}\prod_{k=0}^{i-1}\tfrac{\lambda_{k}}{\mu_{k+1}}}{1+\sum_{n=0}^{\infty}\prod_{k=0}^{n}\tfrac{\lambda_{k}}{\mu_{k+1}}}$
$\leq\frac{C\sum_{i=0}^{\infty}\prod_{k=0}^{i}\tfrac{\lambda_{k}}{\mu_{k+1}}}{1+\sum_{n=0}^{\infty}\prod_{k=0}^{n}\tfrac{\lambda_{k}}{\mu_{k+1}}}$
$<C.$
@@
Proof of Theorem 4.1. We compute the revenue rate of $\pi^{\tilde{\lambda},C-1}$ as
@@
$\mathcal{Z}(\pi^{\tilde{\lambda},\,C-1})$ $=\mathcal{R}(\pi^{\tilde{\lambda},\,C-1})-\mathcal{C}(\pi^{\tilde{\lambda},\,C-1})$
$\geq\mathcal{Z}(\pi^{*})\bigl(1-\mathbb{P}_{C}(\pi^{\tilde{\lambda},\,C-1})\bigr)$
$\Halmos\displaystyle\geq\mathcal{Z}(\pi^{*})\bigl(1-\mathbb{P}_{C}(\pi^{C,\,C-1})\bigr).\quad\Halmos$
The following Lemma 4.6 proves the claimed revenue guarantees of Theorem 4.2.
@@
The following Lemma 4.8 proves the claimed cost guarantee of Theorem 4.2.
@@
$g(\gamma,C)\mathcal{C}(\pi^{*})\geq\mathcal{C}(\pi^{\tilde{\lambda},\gamma})$ (14)
@@
[…]
@@ 5 Sojourn Time Penalty
where the second equality uses Little’s Law. Let $\pi^{*}_{s}$ denote the optimal policy for $\mathcal{Z}_{s}(\cdot)$ which has arrival rates $\lambda^{*}_{s,i}$. We use the notation $\tilde{\lambda}_{s}$ to denote the expected arrival rate under $\pi^{*}_{s}$, i.e, $\tilde{\lambda}_{s}=\sum_{i=0}^{\infty}\lambda^{*}_{s,i}\mathbb{P}_{i}(\pi^{*}_{s})$. In this case, it is no longer possible to obtain guarantees on $\mathcal{Z}_{s}(\cdot)$ using static policies which pick the arrival rate $\tilde{\lambda}_{s}$. This fact is verified with an instance from the class described in Section 3.4. If we take an instance with linear demand ($\lambda=b-ap$) with $b=6000$ and $a=5000$, we can observe that $\mathcal{Z}_{s}(\pi^{*}_{s})\approx 0.17>0$, whereas $\mathcal{Z}_{s}(\pi^{\tilde{\lambda}_{s},0})\approx-0.4<0$. For these instances, taking a cutoff higher than 0 degrades performance further. Thus no result of the flavor of Theorem 4.1 is possible, and hence we must resort to providing bi-criteria bounds on revenue and congestion in the spirit of Theorem 4.2.
In fact, the situation here is worse still: guarantees for the class of static cutoff policies in general are not possible, not just for those which pick $\tilde{\lambda}$. Though for this class we can always attain a non-negative objective value ratio by taking the static policy which picks arrival rates of 0, the performance can be arbitrarily bad. For example, for a single-server system with exponential demand, where $a=.463$ and $b=2$, the optimal static cutoff policy only attains $1\%$ of the optimal objective value.
@@
$E[W(\pi^{\tilde{\lambda}_{s},\gamma})]$ $\leq E[W(\pi^{*}_{s})]\,E[W(\pi^{\tilde{\lambda}_{s},\gamma})]$
$\leq E[W(\pi^{*}_{s})]\,E[W(\pi^{C,\gamma})]$
$=E[W(\pi^{*}_{s})]\cdot\frac{\sum_{i=1}^{\gamma+1}i\,\mathbb{P}_{i}(\pi^{C,\gamma})}{C\bigl(1-\mathbb{P}_{\gamma+1}(\pi^{C,\gamma})\bigr)}.$
@@ 6 Numerical Experiments
In this section, we present the results of numerical experiments on our static policies. Specifically, we consider linear ($\lambda=b-ap$), exponential ($\lambda=be^{-ap}$), and logistic ($\lambda=\frac{b\,(1+e^{-ap_{0}})}{1+e^{a(p-p_{0})}}$) demand functions. We perform experiments with $1,3,5,$ and 1010‐server systems. Though these systems may be relatively small in the context of applications like food delivery and cloud computing, we will see that already with $C=10$ the performance is near-optimal, so further experiments would not be illustrative. For each number of servers $C\in\{1,3,5,10\}$, we sample $a$ uniformly from $[0.1,5]$, $b$ uniformly from $[.5,10]$, and $p_{0}$ uniformly from $[0,20]$. Under linear demand, when $a>b$, the optimal policy does not sell and thus achieves an objective value of 0. Though this is also attainable by a static policy, for simplicity we reject these instances and resample.
@@
    Truncation: For an MM‐state queue, we choose $M$ large enough so that the stationary probability of being at state $M$ is below a small threshold (e.g., 1%). This ensures we capture essentially all of the probability mass.
@@ 7 Conclusion
There are many possible avenues to extend our results. One promising direction of interest would be to relax the assumption that the service times are exponential. With general i.i.d. service times, the optimal pricing policy now depends not only on the number in system but also on the remaining service times of the customers in service. There are many subtleties that must be accounted for to prove similar results, but at the heart of our argument are two very robust tools: Little’s Law and Jensen’s inequality. One may also attempt to provide guarantees that are instance-dependent, using demand and service rate information to provide tighter inequalities. For instance, Eq. (8) uses the simple fact that the total time in the system is at least the average service time, which can be loose in congested settings. Another interesting direction is to consider static pricing in queueing networks that represent vehicle sharing systems, where fair policies are of general concern (elmachtoub2024fair).